OpenStudy (anonymous):

Help please! The perimeter of a rectangular tennis court is 228 meters. The width is 3 meters less than half the length. Find the width and length.

OpenStudy (anonymous):

Length = L, Width = W Perimeter = 2(L) + 2(W) = 228 L/2 -3 = W just solve

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

can i have a medal?

OpenStudy (anonymous):

o i got one :)